# Improving Data Efficiency for LLM Reinforcement Fine-tuning Through Difficulty-targeted Online Data Selection and Rollout Replay

### v1

@@ Abstract
Reinforcement learning (RL) has become an effective approach for fine-tuning large language models (LLMs), particularly to enhance their reasoning capabilities. However, RL fine-tuning remains highly resource-intensive, and existing work has largely overlooked the problem of data efficiency. In this paper, we propose two techniques to improve data efficiency in LLM RL fine-tuning: difficulty-targeted online data selection and rollout replay. We introduce the notion of adaptive difficulty to guide online data selection, prioritizing questions of moderate difficulty that are more likely to yield informative learning signals. To estimate adaptive difficulty efficiently, we develop an attention-based framework that requires rollouts for only a small reference set of questions. The adaptive difficulty of the remaining questions is then estimated based on their similarity to this set. To further reduce rollout cost, we introduce a rollout replay mechanism that reuses recent rollouts, lowering per-step computation while maintaining stable updates. Extensive experiments across $6$ LLM-dataset combinations show that our method reduces RL fine-tuning time by $25\%$ to $65\%$ to reach the same level of performance as the original GRPO algorithm.
@@ 1 Introduction
To further reduce the cost of rollout generation, we introduce a simple rollout replay mechanism. At each training step, we generate a smaller number of new rollouts and reuse past rollouts from recent steps. A bounded First-In-First-Out (FIFO) buffer is maintained to store recent rollouts, from which we retrieve samples to complete each training batch. Although this makes GRPO slightly off-policy, our modified GRPO loss ensures stability, and RR effectively reduces per-step training time without degrading model performance.
@@
  • Extensive experiments on six LLM–dataset combinations show that our method reduces RL fine-tuning time by $25\%$ to $65\%$ while achieving the same performance as the original GRPO algorithm.
Figure 1: Overview of our framework combining difficulty-targeted online data selection and rollout replay. At each training step, the online data selection module selects training questions with adaptive difficulty near 0.5, requiring rollouts on only a small reference set (§4.1, §4.2). The rollout replay module combines current rollouts with retrieved recent rollouts from a FIFO buffer, and the current rollouts are stored into the buffer for future use (§4.3).
@@ 2 Related Work
Data selection seeks to accelerate training by focusing computation on the most informative examples [2]. A key limitation of static data selection methods is their assumption that the importance of samples remains fixed throughout training. Online methods instead periodically reselect data during training to reflect the model’s evolving state [43, 49, 26, 16, 18]. Such adaptability is particularly important in RL, where non-stationary policy updates and environmental interactions necessitate continuous re-evaluation of data utility [31, 44, 33].
@@
On-policy algorithms such as Proximal Policy Optimization (PPO) [38] and Group Relative Policy Optimization (GRPO) [39] have become standard in online RL for LLM reasoning tasks [10]. However, their reliance on freshly collected rollouts for each policy update leads to substantial data inefficiency and computational overhead [24, 4]. Experience replay mitigates this by maintaining a fixed-size buffer of recent transitions collected by the policy. Instead of discarding data after a single use, the buffer enables multiple passes over past rollouts, thereby improving sample efficiency and stabilizing training [6, 51, 36].
@@ 3 Problem Setup
We focus on GRPO algorithm [39] with verifiable rewards. For each question $q$, a group of $G$ individual responses $\left\{o_{i}\right\}_{i=1}^{G}$ are sampled from the old policy $\pi_{\text{old}}$. The advantage of the $i$-th response is calculated by normalizing the group-level rewards $\left\{r_{i}\right\}_{i=1}^{G}$:
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\mathbb{E}_{q\sim\mathcal{D},\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}$
@@ 4 Method
A key challenge in using adaptive difficulty is that computing it requires executing multiple rollouts to obtain rewards, which is one of the most expensive components in LLM RL fine-tuning [^2] For instance, generating rollouts for a batch of 512 samples with maximum sequence length 3072 takes 109.83 seconds on 8 L40S GPUs, nearly half of the total step time.. This raises the question: can we estimate adaptive difficulty efficiently without generating rollouts for all questions? To address this, we propose a lightweight attention-based adaptive difficulty prediction framework that generates rollouts for only a small reference subset of questions. The adaptive difficulty of the remaining questions is then estimated by comparing them to reference questions with known difficulty values using similarity-based attention, thereby avoiding full rollouts. See Fig. 2 for an illustration.
Figure 2: Illustration of our attention-based adaptive difficulty prediction framework. For each unlabeled question, we compute its embedding and attend to reference questions with known difficulty. The predicted difficulty is obtained by computing a similarity-weighted average over the reference difficulties. In this example, the unlabeled question involves inverse trigonometric functions. The model assigns high attention to a reference question that tests a closely related concept and has a difficulty of 1.0. As a result, the predicted difficulty is also close to 1.0. All difficulty values shown correspond to adaptive difficulty scores computed at the same step.
@@ 4.1 Attention-based Adaptive Difficulty Prediction Framework
For the remaining $N-K$ questions, we aim to estimate their adaptive difficulty without performing rollouts. To this end, we employ a lightweight embedding model $E_{\theta}$ to encode questions and capture similarity. We first compute the embeddings $\{z_{i}=E_{\theta}(q_{i})\}_{i=1}^{K}$ for all reference questions. Then, for each unlabeled question $q$, we compute its embedding $z_{q}=E_{\theta}(q)$ and use similarity-weighted averaging to estimate its adaptive difficulty:
$a_{i}=\frac{\exp(z_{q}^{\top}z_{i}/\sqrt{h})}{\sum_{j=1}^{K}\exp(z_{q}^{\top}z_{j}/\sqrt{h})},\quad\hat{d}^{(t)}_{q}=\sum_{i=1}^{K}a_{i}d^{(t)}_{i},$
where $h$ is the embedding dimension.
@@
$\hat{d}^{(t)}_{q,\text{cal}}=\sigma\left(w^{(t)}\cdot\log\frac{\hat{d}^{(t)}_{q}}{1-\hat{d}^{(t)}_{q}}+b^{(t)}\right),$
where $\sigma(\cdot)$ denotes the sigmoid function. The MLP is optimized using binary cross-entropy loss. Full training details can be found in §5.1 and Appendix D.1.
@@ 4.2 Adaptive Difficulty-targeted Online Data Selection
At each training step, with the adaptive difficulty prediction framework, we now efficiently obtain the adaptive difficulty for all questions in the dataset. Inspired by prior work on goal curriculum in RL [7, 45], we prioritize questions whose predicted difficulty is closest to $0.5$.
This selection strategy selects questions that are neither too easy nor too hard for the current policy, as these are intuitively the most informative for learning. Moreover, in GRPO, when all sampled rewards for a question are either 00 or $1$, the group-normalized advantage becomes identically zero, resulting in no gradient signal. By focusing on questions with predicted difficulty near $0.5$, we avoid such degenerate cases and ensure each update contributes meaningfully to policy gradients, thereby accelerating optimization convergence. We formalize this intuition in Theorem 1, which shows that the expected gradient magnitude is maximized when the reward success rate is $0.5$ (i.e., the adaptive difficulty is also $0.5$). A complete proof is provided in Appendix C.
@@
Consider a single question $q$, where $G$ responses $\{o_{i}\}_{i=1}^{G}$ are sampled independently from the current policy $\pi_{\theta}(\cdot\mid q)$. Each response receives a binary reward $r_{i}\in\{0,1\}$, sampled i.i.d. from a Bernoulli$(p)$ distribution, where $p$ represents the reward success rate. Define the group-relative advantage $\hat{A}_{i}$ as in Eq. 1. We consider the unclipped policy gradient estimator for this question without KL penalty:
$g=\sum_{i=1}^{G}\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q).$
Assume that the likelihood gradients $\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)$ are independent of the reward distribution parameter $p$ and have bounded variance. Then, the expected squared norm of the gradient satisfies:
@@
1:Initial policy model $\pi_{\theta}$, reward model $r_{\varphi}$, math question dataset $\mathcal{D}$, target difficulty $\alpha$, batch size $B$, total steps $T$, reference set size $K$, sampling temperature $\tau$, adaptive difficulty prediction framework DiffPred (§4.1), fresh rollout fraction $\delta\in(0,1]$, buffer capacity $C$
@@
3:for step $=1,\dots,T$ do
4:       // Adaptive difficulty prediction
5:     Sample reference set $\mathcal{D}_{\text{ref}}\subset\mathcal{D}$ uniformly at random, where $|\mathcal{D}_{\text{ref}}|=K$
6:     for each $q\in\mathcal{D}_{\text{ref}}$ do
7:         Generate $G$ outputs $\{o^{q}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$
8:         Compute rewards $r_{i}^{q}=r_{\varphi}(o_{i}^{q})$ for $i\in[G]$ and difficulty score $d_{q}=\frac{1}{G}\sum_{i=1}^{G}(1-r_{i}^{q})$
9:     end for
10:     Predict adaptive difficulty $\hat{d}_{q^{\prime}}=\textsc{DiffPred}(\mathcal{D}_{\text{ref}},\{d_{q}\},q^{\prime})$ for all $q^{\prime}\in\mathcal{D}\setminus\mathcal{D}_{\text{ref}}$.
11:     Sample rollout batch $\mathcal{B}_{\text{rollout}}$ of size $\delta B$ from $\mathcal{D}$ according to:
@@
12:     for each $q\in\mathcal{B}_{\text{rollout}}$ do
13:         Generate $G$ outputs $\{o^{q}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$
14:         Compute rewards $r_{i}^{q}=r_{\varphi}(o_{i}^{q})$ for $i\in[G]$ and group average reward $\bar{r}_{q}=\frac{1}{G}\sum_{i=1}^{G}r_{i}^{q}$
15:         Obtain advantages $\hat{A}_{i}^{q}$ and policy probabilities $\pi_{\theta_{\text{old}}}(o_{i}^{q}\mid q)$ for $i\in[G]$
16:     end for
17:       // Rollout Replay and Update
@@
18:     Sample $(1-\delta)B$ samples from buffer $\mathcal{R}$ to complete batch $\mathcal{B}$
@@
19:     Update policy $\pi_{\theta}$ using modified GRPO objective on batch $\mathcal{B}$
20:       // Store informative rollouts in buffer
21:     Add $(q,o_{i}^{q},\hat{A}_{i}^{q},\pi_{\theta_{\text{old}}}(o_{i}^{q}\mid q))$ to $\mathcal{R}$ for $q\in\mathcal{B}_{\text{rollout}},\ \bar{r}_{q}\notin\{0,1\},\ i\in[G]$
22:     Remove the oldest samples from $\mathcal{R}$ if buffer is full until $|\mathcal{R}|\leq C$
23:end for
4.3 Rollout Replay
@@
To further improve data efficiency, we aim to reduce the time cost of each training step. Since rollout generation is one of the most expensive components, we adopt a rollout replay mechanism. Specifically, at each training step, we generate new rollouts for only a fraction $\delta B$ of the batch, where $\delta\in(0,1]$, and fill the remaining $(1-\delta)B$ samples using recent rollouts sampled from a FIFO replay buffer $\mathcal{D}_{\text{replay}}$ with capacity $C$.
@@
However, naively reusing past rollouts introduces bias into the policy gradient estimation, as the data is no longer drawn from the current policy. This mismatch can lead to unstable training and performance degradation [28]. Inspired by off-policy variants of PPO [35], we propose a modified GRPO loss using importance sampling with respect to the behavior policy $\pi_{\text{behavior}}$ under which the rollouts stored in the buffer were originally collected:
@@
$\mathcal{J}_{\text{GRPO-RR}}(\theta)=$ $\mathbb{E}_{q\sim\mathcal{D}_{\text{replay}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\text{behavior}}(\cdot\mid q)}}$
@@
$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left(\min\left(\tilde{r}_{i,t}(\theta)A_{i},\ \text{clip}(\tilde{r}_{i,t}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\right)\Bigg{]},$
where $\tilde{r}_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\text{behavior}}(o_{i,t}|q,o_{i,<t})}$. By appropriately controlling the buffer size $C$, we empirically show that rollout replay improves sample efficiency while maintaining training stability.
@@
For each newly generated rollout, if the group average reward is neither 0 nor 1 (i.e., the sample yields a non-zero gradient signal), we store the question, its sampled rollouts, computed advantages, and policy probabilities into the buffer. When the buffer is full, the oldest samples are discarded.

### v2

@@ Abstract
Reinforcement learning (RL) has become an effective approach for fine-tuning large language models (LLMs), particularly to enhance their reasoning capabilities. However, RL fine-tuning remains highly resource-intensive, and existing work has largely overlooked the problem of data efficiency. In this paper, we propose two techniques to improve data efficiency in LLM RL fine-tuning: difficulty-targeted online data selection and rollout replay. We introduce the notion of adaptive difficulty to guide online data selection, prioritizing questions of moderate difficulty that are more likely to yield informative learning signals. To estimate adaptive difficulty efficiently, we develop an attention-based framework that requires rollouts for only a small reference set of questions. The adaptive difficulty of the remaining questions is then estimated based on their similarity to this set. To further reduce rollout cost, we introduce a rollout replay mechanism inspired by experience replay in traditional RL. This technique reuses recent rollouts, lowering per-step computation while maintaining stable updates. Experiments across $6$ LLM-dataset combinations show that our method reduces RL fine-tuning time by $23\%$ to $62\%$ while reaching the same level of performance as the original GRPO algorithm. Our code repository is available at https://github.com/ASTRAL-Group/data-efficient-llm-rl/.
@@ 1 Introduction
To further reduce the cost of rollout generation, we introduce a simple rollout replay mechanism, motivated by experience replay in standard RL [6]. At each training step, we generate fewer new rollouts and reuse past rollouts from recent steps. A bounded First-In-First-Out (FIFO) buffer is maintained to store recent rollouts, from which we retrieve samples to complete each training batch. Although this makes GRPO slightly off-policy, our modified GRPO loss ensures stability, and thus RR effectively reduces per-step training time without degrading model performance.
Figure 1: Overview of our framework combining Difficulty-targeted Online Data Selection and Rollout Replay. At each training step, the online data selection module selects training questions with adaptive difficulty near 0.5, requiring rollouts only on a small reference set (§4.1, §4.2). The rollout replay module combines current rollouts with retrieved recent rollouts from a FIFO buffer, and the current rollouts are stored into the buffer for future use (§4.3).
@@
  • Extensive experiments on six LLM–dataset combinations show that our method reduces RL fine-tuning time by $23\%$ to $62\%$ while achieving the same performance as the original GRPO algorithm.
@@ 2 Related Work
Data selection seeks to accelerate training by focusing computation on the most informative examples [2]. A key limitation of static data selection methods is their assumption that the importance of samples remains fixed throughout training. Online methods instead periodically reselect data during training to reflect the model’s evolving state [43, 49, 26, 16, 18]. Such adaptability is particularly important in RL, where non-stationary policy updates and environment dynamics necessitate continuous re-evaluation of data utility [31, 44, 33].
@@
On-policy algorithms such as Proximal Policy Optimization (PPO) [38] and Group Relative Policy Optimization (GRPO) [39] have become standard choices for online RL fine-tuning in LLM reasoning tasks [9]. However, their reliance on freshly collected rollouts for each policy update leads to substantial data inefficiency and computational overhead [24, corrado2023policy]. Experience replay mitigates this by maintaining a fixed-size buffer of recent transitions collected by the policy. Instead of discarding data after a single use, the buffer enables multiple passes over past rollouts, thereby improving sample efficiency and stabilizing training [6, 51, 36].
@@ 3 Problem Setup
We focus on the GRPO algorithm [39] with verifiable rewards. For each question $q$, a group of $G$ individual responses $\left\{o_{i}\right\}_{i=1}^{G}$ are sampled from the old policy $\pi_{\text{old}}$. The advantage of the $i$-th response is calculated by normalizing the group-level rewards $\left\{r_{i}\right\}_{i=1}^{G}$, where $r_{i}\in\{0,1\}$:
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\$$\mathbb{E}_{q\sim\mathcal{D},\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}$
@@ 4 Method
A key challenge in using adaptive difficulty is that computing it requires executing multiple rollouts, which is one of the most expensive components in LLM RL fine-tuning [^2] For instance, generating rollouts for a batch of 512 samples with maximum sequence length 3072 takes 109.83 seconds on 8 L40S GPUs, nearly half of the total step time.. This raises the question: can we estimate adaptive difficulty efficiently without generating rollouts for all questions? To address this, we propose a lightweight attention-based adaptive difficulty prediction framework that generates rollouts for only a small reference subset of questions. The adaptive difficulty of the remaining questions is then estimated by comparing them to reference questions with known difficulty values using similarity-based attention, thereby avoiding full rollouts. See Fig. 2 for an illustration.
Figure 2: Illustration of our attention-based adaptive difficulty prediction framework. For each unlabeled question, we compute its embedding and attend to reference questions to obtain similarity scores. The predicted difficulty of the unlabeled question is obtained by computing an attention-weighted average, where similarities to reference questions serve as attention scores over their associated difficulties. In this example, the unlabeled question involves inverse trigonometric functions. The model assigns high attention to a reference question that tests a closely related concept and has a difficulty of 1.0. As a result, the predicted difficulty is also close to 1.0. All difficulty values shown correspond to adaptive difficulty scores computed at the same step.
@@ 4.1 Attention-based Adaptive Difficulty Prediction Framework
For the remaining NKN-K questions, we aim to estimate their adaptive difficulty without performing rollouts. To this end, we employ a lightweight embedding model $E_{\theta}$ to encode questions and capture similarity. We first compute the embeddings $\{z_{i}=E_{\theta}(q_{i})\}_{i=1}^{K}$ for all reference questions. Denote $h$ as the embedding dimension. Then, for each unlabeled question $q$, we compute its embedding $z_{q}=E_{\theta}(q)$ and use similarity-weighted averaging to estimate its adaptive difficulty:
$a_{i}=\frac{\exp(z_{q}^{\top}z_{i}/\sqrt{h})}{\sum_{j=1}^{K}\exp(z_{q}^{\top}z_{j}/\sqrt{h})},\quad\hat{d}^{(t)}_{q}=\sum_{i=1}^{K}a_{i}d^{(t)}_{i}$.
@@
$\hat{d}^{(t)}_{q,\text{cal}}=\sigma\left(w^{(t)}\cdot\left(\log\hat{d}^{(t)}_{q}-\log(1-\hat{d}^{(t)}_{q})\right)+b^{(t)}\right),$
where $\sigma(\cdot)$ denotes the sigmoid function. The MLP is optimized using binary cross-entropy loss. Full training details can be found in §5.1 and Appendix C.1.
@@ 4.2 Adaptive Difficulty-targeted Online Data Selection
At each training step, with the adaptive difficulty prediction framework, we now efficiently obtain the adaptive difficulty for all questions in the training set. Inspired by prior work on goal curriculum in RL [7, 45], we prioritize questions whose predicted difficulty is closest to $0.5$.
This selection strategy selects questions that are neither too easy nor too hard for the current policy, as these are intuitively the most informative for learning. Moreover, in GRPO, when all sampled rewards for a question are either 0 or $1$, the group-normalized advantage becomes identically zero, resulting in no gradient signal. By focusing on questions with predicted difficulty near $0.5$, we avoid such degenerate cases and ensure each update contributes meaningfully to policy gradients, thereby accelerating optimization convergence. We formalize this intuition in Theorem 1, which shows that the expected gradient magnitude is maximized when the reward success rate is $0.5$ (i.e., the adaptive difficulty is also $0.5$). A complete proof is provided in Appendix B.
@@
Consider a single question $q$, where $G$ responses $\{o_{i}\}_{i=1}^{G}$ are sampled independently from the current policy $\pi_{\theta}(\cdot\mid q)$. Each response receives a binary reward $r_{i}\in\{0,1\}$, sampled i.i.d. from a Bernoulli(p)(p) distribution, where $p$ represents the reward success rate. Define the group-relative advantage $\hat{A}_{i}$ as in Eq. 1. We consider the unclipped policy gradient estimator for this question without KL penalty $g=\sum_{i=1}^{G}\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q).$ Under mild assumptions on the reward and the likelihood gradients $\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)$ (detailed in Appendix B), the expected squared norm of the gradient satisfies:
@@
Discussion: Our dynamic selection mechanism implicitly promotes diversity.
As questions near the target difficulty are repeatedly selected and trained on, their predicted difficulty gradually deviates from 0.5. They are then less likely to be sampled again, allowing other under-explored questions to enter the selection pool. This dynamic prevents overfitting to a small subset of questions and encourages broader coverage over time.
4.3 Rollout Replay
@@
1:Initial policy model $\pi_{\theta}$, reward model $r_{\varphi}$, training dataset $\mathcal{D}$, target difficulty $\alpha$, batch size $B$, total steps $T$, reference set size $K$, sampling temperature $\tau$, adaptive difficulty prediction framework DiffPred (§4.1), fresh rollout fraction $\delta\in(0,1]$, buffer capacity $C$
@@
3:Set $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
4:for step $1,\dots,T$ do
5:    // Adaptive difficulty prediction
6:  Sample reference set $\mathcal{D}_{\text{ref}}\subset\mathcal{D}$ uniformly at random, where $|\mathcal{D}_{\text{ref}}|=K$
7:  for each $q\in\mathcal{D}_{\text{ref}}$ do
8:   Generate $G$ outputs $\{o^{q}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$
9:   Compute rewards $r_{i}^{q}=r_{\varphi}(o_{i}^{q})$ for $i\in[G]$ and difficulty score $d_{q}=\frac{1}{G}\sum_{i=1}^{G}(1-r_{i}^{q})$
10:  end for
11:  Predict adaptive difficulty $\hat{d}_{q^{\prime}}=\textsc{DiffPred}(\mathcal{D}_{\text{ref}},\{d_{q}\mid q\in\mathcal{D}_{\text{ref}}\},q^{\prime})$ for all $q^{\prime}\in\mathcal{D}\setminus\mathcal{D}_{\text{ref}}$
12:  Sample rollout batch $\mathcal{B}_{\text{rollout}}$ of size $\delta B$ from $\mathcal{D}$ according to:
@@
13:  for each $q\in\mathcal{B}_{\text{rollout}}$ do
14:   Generate $G$ outputs $\{o^{q}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$
15:   Compute rewards $r_{i}^{q}=r_{\varphi}(o_{i}^{q})$ for $i\in[G]$ and group average reward $\bar{r}_{q}=\frac{1}{G}\sum_{i=1}^{G}r_{i}^{q}$
16:   Obtain advantages $\hat{A}_{i}^{q}$ and policy probabilities $\pi_{\theta_{\text{old}}}(o_{i}^{q}\mid q)$ for $i\in[G]$
17:  end for
18:   // Rollout Replay and update
@@
19:  Sample $(1-\delta)B$ samples from buffer $\mathcal{R}$ to complete batch $\mathcal{B}$ if $|\mathcal{R}|\geq(1-\delta)B$
@@
20:  Update policy $\pi_{\theta}$ using modified GRPO objective $\mathcal{J}_{\text{GRPO-RR}}$ on batch $\mathcal{B}$
21:   // Store informative rollouts in buffer
22:  Add $\bigl(q,\;\{\,(o_{i}^{q},\;\hat{A}_{i}^{q},\;\pi_{\theta_{\text{old}}}(o_{i}^{q}\mid q))\,\}_{i=1}^{G}\bigr)$ to $\mathcal{R}$ for $q\in\mathcal{B}_{\text{rollout}},\ \bar{r}_{q}\notin\{0,1\}$
23:  Remove the oldest samples from $\mathcal{R}$ until $|\mathcal{R}|\leq C$
24:  Set $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
25:end for
@@
To further improve data efficiency, we aim to reduce the time cost of each training step. Since rollout generation is one of the most expensive components, we adopt a rollout replay mechanism, inspired by experience replay in traditional RL. Specifically, at each training step, we generate new rollouts for only a fraction $\delta B$ of the batch, where $\delta\in(0,1]$, and fill the remaining $(1-\delta)B$ samples using recent rollouts sampled from a FIFO replay buffer $\mathcal{D}_{\text{replay}}$ with capacity $C$.
@@
However, naively reusing past rollouts introduces bias into the policy gradient estimation, as the data is no longer drawn from the current policy. This mismatch can lead to unstable training and performance degradation [28]. Inspired by off-policy variants of PPO [35], we propose a modified GRPO loss using importance sampling with respect to the behavior policy $\pi_{\theta_{\text{behavior}}}$ under which the rollouts stored in the buffer were originally collected:
@@
$\mathcal{J}_{\text{GRPO-RR}}(\theta)=\$$\mathbb{E}_{q\sim\mathcal{D},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{behavior}}}(\cdot\mid q)}}$
@@
$\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left(\min\left(\tilde{r}_{i,t}(\theta)A_{i},\ \text{clip}(\tilde{r}_{i,t}(\theta),1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\right)\Bigg],\vskip-14.22636pt$
@@
where $\tilde{r}_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{behavior}}}(o_{i,t}|q,o_{i,<t})}$. By appropriately controlling the buffer size $C$, we empirically demonstrate that rollout replay improves sample efficiency while maintaining training stability. For each newly generated rollout, if the group average reward is neither 0 nor 1 (i.e., the sample yields a non-zero gradient signal), we store the question, its sampled rollouts, computed advantages, and policy probabilities into the buffer. When the buffer is full, the oldest samples are discarded.